OpenStudy (anonymous):

The length of a certain rectangle is six more than three times its width. If the width of the rectangle is 4 units, what is its length?

OpenStudy (texaschic101):

L = 3W + 6 W = 4 L = 3(4) + 6 L = 12 + 6 L = 18

OpenStudy (anonymous):

Thank you!!!

OpenStudy (texaschic101):

sure thing :)

OpenStudy (anonymous):

wana help me with another one :/

OpenStudy (anonymous):

The length of a rectangle is 2 centimeters longer than its width. The perimeter is 16 centimeters. What are the length and width of the rectangle?